A 50-kg woman cheats on her diet and eats a 575-Calorie (575 kcal) jelly doughnut for breakfast.

(a) How many joules of energy are the equivalent of one jelly doughnut?
J

(b) How many stairs must the woman climb to perform an amount of mechanical work equivalent to the food energy in one jelly doughnut? Assume the height of a single stair is 15 cm.
stairs

(c) If the human body is only 26% efficient in converting chemical energy to mechanical energy, how many stairs must the woman climb to work off her breakfast?
stairs
Answer #1

1 calorie = 4.186 joules

So

Doughnut has 2406.95 joules

Let number of stairs = n

Increase in potential energy = n x mgh = n × 73.5

So n = 2406.95/73.5 = 33

If efficiency is 26 %

Then n = (0.26 × 2406.95)/73.5 = 9

Please rate it up thanks :)